you invested $29,000 in two accounts paying 8% and 9% annual interest respectively. if the total interest earned for the year was $2350 how much was invested at each rate?
Answer by Maths68(1474)


Amount invested at 8% = x
Amount invested at 9% = 29000-x
Given
Total interest on both investments = 2350
8% * (x) + 9% * (29000-x) = 2350
8/100 * (x) + 9/100 * (29000-x) = 2350
8/100 * (x) + 9/100 * (29000-x) = 2350
Multiply by 100 both sides
8 * (x) + 9 * (29000-x) = 2350 * 100
8x + 261000-9x = 235000
-x=235000-261000
-x=-26000
Multiply by -1 boths sides
x=26000
Amount invested at 8% = x =26000
Amount invested at 9% =29000-x=29000-26000 = 3000
Check
=====
26000 * 8/100 +3000 * 9/100 = 2350
260*8+30*9=2350
2080+270=2350
2350=2350